OpenStudy (anonymous):

please help me solve x^4 y^17 / y^29 x^10 and answer must be written with positive exponet

OpenStudy (anonymous):

\[\frac{ A^x }{ A^y }=A^{x-y}|A \ne 0\]Can you use this exponent law to figure it out?

OpenStudy (anonymous):

Write it as 2 fractions. Like this:\[\frac{ x^4 y^{17}}{ x^{10} y^{29} }=\frac{ x^4 }{ x^{10} }*\frac{ y^{17} }{ y^{29} }\]And then use the above exponent law to evaluate. @nikki2

OpenStudy (anonymous):

I am not too sure about the above law. How would I write the X & Y in the format you mentioned aboue

OpenStudy (anonymous):

i think i figured it out would i just use the A as the base and the x and y as subscript rasied to their respective powers and do the computation to get the positive exponent